# BGE M3-Embedding: Multi-Lingual, Multi-Functionality, Multi-Granularity Text Embeddings Through Self-Knowledge Distillation

### v3

@@ 4 Experiment
In the following part, we evaluate our model on three tasks: multi-lingual retrieval, cross-lingual retrieval, and long-doc retrieval.
@@
4.1 Multi-Lingual Retrieval
We evaluate the multi-lingual retrieval performance with MIRACL Zhang et al. (2023c), which consists of ad-hoc retrieval tasks in 18 languages. Each task is made up of query and passage presented in the same language. Following the official benchmark, we evaluate our method using Pyserini Lin et al. (2021), and use nDCG@10 as the primary evaluation metric (Recall@100 is also measured and reported in Appendix C.1). We incorporate the following baselines in our experiment: the lexical retrieval method: BM25 Robertson and Zaragoza (2009); the dense retrieval methods: mDPR333https://huggingface.co/castorini/mdpr-tied-pft-msmarco Zhang et al. (2023b), mContriever444https://huggingface.co/facebook/mcontriever-msmarco Izacard et al. (2022), mE5 Wang et al. (2022) and E5 Wang et al. (2023). To make the BM25 and M3 more comparable, in the experiment, we use the same tokenizer as M3 (i.e., the tokenizer of XLM-Roberta) for BM25. Using the same vocabulary from XLM-Roberta can also ensure that both approaches have the same retrieval latency. The results of BM25 with different tokenizers are shown in Appendix C.2. We also make a comparison with Text-Embedding-3-Large(abbreviated as OpenAI-3), which was recently released by OpenAI555https://platform.openai.com/docs/guides/embeddings.
We can make the following observations according to the experiment result in Table 2. Firstly, M3-Embedding already achieves a superior retrieval performance with only its dense retrieval functionality (denoted as Dense). It not only outperforms other baseline methods in the average performance, but also maintains a consistent empirical advantage in most of individual languages. Even compared with E5, which leverages a much larger Mistral-7B model as the text encoder and specifically trained with English data, our method is able to produce a similar result in English and notably higher results in the other languages. Besides, the sparse retrieval functionality (denoted as Sparse) is also effectively trained by M3-Embedding, as it outperforms the typical BM25 methods in all languages. We can also observe the additional improvement from multi-vector retrieval666Mult-vec is used to re-rank the top-200 candidates from Dense for efficient processing. (denoted as Mult-vec), which relies on fine-grained interactions between query and passage’s embeddings to compute the relevance score. Finally, the collaboration of dense and sparse method, e.g., Dense+Sparse777Retrieve the top-1000 candidates with dense and sparse method; then re-rank using the sum of two scores., leads to a further improvement over each individual method; and the collaboration of all three methods888Re-rank based on the sum of all three scores. (denoted as All) brings forth the best performance.
@@ 4.1 Multi-Lingual Retrieval
| text-embedding-3-large | 8192 | 51.6 |
@@
4.2 Cross-Lingual Retrieval
We make evaluation for the cross-lingual retrieval performance with the MKQA benchmark Longpre et al. (2021), which includes queries in 25 non-English languages. For each query, it needs to retrieve the ground-truth passage from the English Wikipedia corpus. In our experiment, we make use of the well-processed corpus offered by the BEIR999https://huggingface.co/datasets/BeIR/nq Thakur et al. (2021). Following the previous study Karpukhin et al. (2020), we report Recall@100 as the primary metric (Recall@20 is reported as an auxiliary metric in the Appendix).
4.3 Multilingual Long-Doc Retrieval
We evaluate the retrieval performance with longer sequences with two benchmarks: MLDR (Multilingual Long-Doc Retrieval), which is curated by the multilingual articles from Wikipedia, Wudao and mC4 (see Table 8), and NarrativeQA101010Using the evaluation pipeline from Günther et al. (2024) s Koˇciský et al. (2018); Günther et al. (2024), which is only for English. In addition to the previous baselines, we further introduce JinaEmbeddingv2111111https://huggingface.co/jinaai/jina-embeddings-v2-base-en Günther et al. (2024), text-embedding-ada-002 and text-embedding-3-large from OpenAI given their outstanding long-doc retrieval capability.
| M3-w.skd | |||||||||||||||||||
| Dense | 67.8 | 78.4 | 80.0 | 56.9 | 55.5 | 57.7 | 78.6 | 57.8 | 59.3 | 56.0 | 72.8 | 69.9 | 70.1 | 78.6 | 86.2 | 82.6 | 61.7 | 56.8 | 60.7 |
| Sparse | 53.9 | 67.1 | 68.7 | 43.7 | 38.8 | 45.2 | 65.3 | 35.5 | 48.2 | 48.9 | 56.3 | 61.5 | 44.5 | 57.9 | 79.0 | 70.9 | 36.3 | 32.2 | 70.0 |
| Multi-vec | 69.0 | 79.6 | 81.1 | 59.4 | 57.2 | 58.8 | 80.1 | 59.0 | 61.4 | 58.2 | 74.5 | 71.2 | 71.2 | 79.0 | 87.9 | 83.0 | 62.7 | 57.9 | 60.4 |
| M3-w.o.skd | |||||||||||||||||||
| Dense | 67.2 | 78.0 | 79.1 | 56.4 | 54.9 | 57.4 | 78.3 | 57.8 | 58.9 | 55.1 | 72.3 | 68.7 | 69.5 | 77.8 | 85.8 | 82.5 | 62.1 | 55.9 | 59.9 |
| Sparse | 36.7 | 48.2 | 52.1 | 24.3 | 20.3 | 25.9 | 48.6 | 16.9 | 30.1 | 32.1 | 33.0 | 43.1 | 27.1 | 45.3 | 63.8 | 52.0 | 22.6 | 16.5 | 59.4 |
| Multi-vec | 67.8 | 78.7 | 80.2 | 57.6 | 56.1 | 57.4 | 79.0 | 57.9 | 59.2 | 57.5 | 74.0 | 70.3 | 70.2 | 78.6 | 86.9 | 82.1 | 61.0 | 56.7 | 57.4 |
@@ 4.4 Ablation study
Self-knowledge distillation This ablation study is performed to analyze the impact of self-knowledge distillation (skd). Particularly, we disable the distillation processing and have each retrieval method trained independently (denoted as M3-w.o.skd). According to our evaluation on MIRACL (Table 6), the original method, i.e. M3 w.skd, brings in better performances than the ablation method in all settings, i.e., Dense, Sparse, Multi-vec. Notably, the impact is more pronounced for sparse retrieval, which indicates the incompatibility between dense and sparse retrieval methods.
Impact of multi-stage training We also conducted experiments to explore the impact of different stages. Fine-tuning indicates fine-tuning directly from the xlm-roberta Conneau et al. (2020) model, and RetroMAE+Fine-tuning refers to fine-tuning on a model trained with RetroMAE Xiao et al. (2022). Meanwhile, RetroMAE+Unsup+Fine-tuning involves fine-tuning on a model trained with RetroMAE and then pre-trained on unsupervised data. The results are summarized in Table 7. We can see that RetroMAE can significantly improve retrieval performance, and pre-training on unsupervised data can further enhance the retrieval ability of the embedding model.
@@
| Fine-tune | 59.3 |
| RetroMAE + Fine-tune | 64.8 |
| RetroMAE + Unsup + Fine-tune | 67.8 |
@@ 5 Conclusion
In this paper, we present M3-Embedding, which achieves notable versatility in supporting multi-lingual retrieval, handling input of diverse granularities, and unifying different retrieval functionalities. We perform comprehensive and high-quality curation of training data, optimize the learning process with self-knowledge distillation, and improve the training through and batch size with efficient batching. The effectiveness of M3-Embedding is verified by our experimental studies, where it leads to superior performances on multi-lingual retrieval, cross-lingual retrieval, and multi-lingual long-document retrieval tasks.
@@ B.1 Experimental Hyperparameters
We adopt a further pre-trained XLM-RoBERTa121212https://huggingface.co/FacebookAI/xlm-roberta-large as the foundational model. We extend the max position to 8192 and update the model via the RetroMAE Xiao et al. (2022) method. The data comprises Pile Gao et al. (2020), Wudao Yuan et al. (2021), and mC4 Raffel et al. (2019) datasets. We sampled a total of 184 million text samples from these sources, covering 105 languages. The maximum sequence length is 8192 and the learning rate is . The batch size is set to 32 and we accumulate the gradient over 16 steps. Pre-training is conducted on 32 A100(40GB) GPUs for 20,000 steps.
@@ C.2 Different Tokenizer for BM25
Using the Analyzer from Lucene131313https://github.com/apache/lucene/tree/main/lucene/analysis/ can significantly enhance the effectiveness of BM25. Lucene analyzer includes multiple steps typically including tokenization, stemming, stopword removal, etc, achieving better results than directly using the tokenzier of xlm-roberta. Additionally, it’s worth noting that the vocabulary size of the tokenizer from xlm-roberta is limited, resulting in fewer unique tokens after encoding documents (for example, on the MLDR dataset, the tokenizer of xlm-roberta produces 1056 unique terms per article, while Lucene’s analyzer generates 1451 unique terms, which is over 37% more and will increase retrieval latency).
@@
| Dense | 93.8 | 97.6 | 98.7 | 90.7 | 90.1 | 89.6 | 97.9 | 93.1 | 94.2 | 90.5 | 97.5 | 95.5 | 95.9 | 97.2 | 99.4 | 99.1 | 95.7 | 90.8 | 74.2 |
| Sparse | 85.6 | 92.0 | 96.6 | 81.5 | 71.9 | 87.0 | 91.5 | 73.2 | 87.2 | 84.9 | 92.3 | 91.8 | 77.0 | 85.1 | 98.0 | 95.2 | 72.8 | 69.0 | 92.9 |
| Multi-vec | 94.5 | 97.8 | 98.9 | 91.6 | 91.3 | 90.5 | 98.2 | 95.4 | 94.9 | 92.5 | 98.0 | 95.9 | 96.5 | 97.3 | 99.4 | 99.2 | 96.2 | 92.3 | 74.6 |
| Dense+Sparse | 94.4 | 98.0 | 98.9 | 92.4 | 91.5 | 91.0 | 98.4 | 93.9 | 95.3 | 92.6 | 97.5 | 95.6 | 96.6 | 97.6 | 99.1 | 99.0 | 95.7 | 90.8 | 75.4 |
| All | 94.6 | 98.0 | 98.9 | 92.1 | 91.8 | 91.2 | 98.4 | 95.0 | 94.9 | 92.4 | 97.9 | 96.0 | 96.7 | 97.2 | 99.4 | 99.2 | 96.5 | 92.2 | 74.6 |

### v4

@@ 4 Experiment
In this section, we investigate M3-Embedding’s performance in terms of multi-lingual retrieval, cross-lingual retrieval, and long-doc retrieval. We also explore the impact of its technical factors.
4.1 Multi-Lingual Retrieval
We evaluate the multi-lingual retrieval performance with MIRACL Zhang et al. (2023c), which consists of ad-hoc retrieval tasks in 18 languages. Each task is made up of query and passage presented in the same language. Following the official benchmark, we evaluate our method using Pyserini Lin et al. (2021), and use nDCG@10 as the primary evaluation metric (Recall@100 is also measured and reported in Appendix C.1). Specifically, for the dense method (denoted as Dense), we first use it to generate the embeddings of the corpus and then build the dense index for searching top-1000 candidates with Faiss. For the sparse method (denoted as Sparse), we first use it to generate the weights of the corpus and then build the sparse index for searching top-1000 candidates with Lucene. For the multi-vector method (denoted as Multi-vec), considering its heavy cost, we use it as reranker to re-rank the top-200 candidates from dense method. For the hybrid retrieval of dense method and sparse method (denoted as Dense+Sparse), we set , and in equation(1) to re-rank the union set of top-1000 candidates from Dense and top-1000 candidate from Sparse. For the hybrid retrieval of all three methods (denoted as All), we set , and in equation(1) to re-rank the top-200 candidates from Dense.
We incorporate the following baselines in our experiment: the lexical retrieval method: BM25 Robertson and Zaragoza (2009); the dense retrieval methods: mDPR333https://huggingface.co/castorini/mdpr-tied-pft-msmarco Zhang et al. (2023b), mContriever444https://huggingface.co/facebook/mcontriever-msmarco Izacard et al. (2022), mE5 Wang et al. (2022) and E5 Wang et al. (2023). To make the BM25 and M3 more comparable, in the experiment, we use the same tokenizer as M3 (i.e., the tokenizer of XLM-Roberta) for BM25. Using the same vocabulary from XLM-Roberta can also ensure that both approaches have the same retrieval latency. The results of BM25 with different tokenizers are shown in Appendix C.2. We also make a comparison with Text-Embedding-3-Large(abbreviated as OpenAI-3), which was recently released by OpenAI555https://platform.openai.com/docs/guides/embeddings.
@@ 4.1 Multi-Lingual Retrieval
We can make the following observations according to the experiment result in Table 1. Firstly, M3-Embedding already achieves a superior retrieval performance with only its dense retrieval functionality (Dense). It not only outperforms other baseline methods in the average performance, but also maintains a consistent empirical advantage in most of individual languages. Even compared with E5, which leverages a much larger Mistral-7B model as the text encoder and specifically trained with English data, our method is able to produce a similar result in English and notably higher results in the other languages. Besides, the sparse retrieval functionality (Sparse) is also effectively trained by M3-Embedding, as it outperforms the typical BM25 methods in all languages. We can also observe the additional improvement from multi-vector retrieval, which relies on fine-grained interactions between query and passage’s embeddings to compute the relevance score. Finally, the collaboration of dense and sparse method (Dense+Sparse) leads to a further improvement over each individual method, and the collaboration of all three methods (All) brings forth the best performance.
@@
4.2 Cross-Lingual Retrieval
We make evaluation for the cross-lingual retrieval performance with the MKQA benchmark Longpre et al. (2021), which includes queries in 25 non-English languages. For each query, it needs to retrieve the passages containing answers from the English Wikipedia corpus. In our experiment, we make use of the well-processed corpus offered by the BEIR666https://huggingface.co/datasets/BeIR/nq Thakur et al. (2021). Following the previous study Izacard et al. (2022), we report Recall@100 as the primary metric (Recall@20 is reported as an auxiliary metric in the Appendix C.1). For Dense+Sparse method and All method, we set the same weights as in MIRACL dataset.
4.3 Multilingual Long-Doc Retrieval
We evaluate the retrieval performance with longer sequences with two benchmarks: MLDR (Multilingual Long-Doc Retrieval), which is curated by the multilingual articles from Wikipedia, Wudao and mC4 (see Table 7), and NarrativeQA Kočiský et al. (2018); Günther et al. (2023), which is only for English. In addition to the previous baselines, we further introduce JinaEmbeddingv2 Günther et al. (2023), text-embedding-ada-002 and text-embedding-3-large from OpenAI given their outstanding long-doc retrieval capability. For Dense+Sparse method, we set , and in equation(1). For All method, we set , and in equation(1).
@@ 4.3 Multilingual Long-Doc Retrieval
| text-embedding-3-large | 8191 | 51.6 |
@@ 4.4 Ablation study
Self-knowledge distillation. The ablation study is performed to analyze the impact of self-knowledge distillation (skd). Particularly, we disable the distillation processing and have each retrieval method trained independently (denoted as M3-w.o.skd). According to our evaluation on MIRACL (Table 5), the original method, i.e. M3-w.skd, is able to achieve better performances than the ablation method in all settings, i.e., Dense, Sparse, Multi-vec. Notably, the impact is more pronounced for sparse retrieval. Such a result also reflects the incompatibility between dense and sparse retrieval methods. With skd, the incompatibility can be largely overcome. (More detailed results are available in Appendix C.1.)
| Model | MIRACL | |
| M3-w.skd | Dense | 69.2 |
| Sparse | 53.9 | |
| Multi-vec | 70.5 | |
| M3-w.o.skd | Dense | 68.7 |
| Sparse | 36.7 | |
| Multi-vec | 69.3 | |
@@
| Fine-tune | 60.5 |
| RetroMAE + Fine-tune | 66.1 |
| RetroMAE + Unsup + Fine-tune | 69.2 |
Impact of multi-stage training. We also make explorations for the impacts from different training stages. Fine-tuning indicates the direct fine-tuning from XLM-RoBERTA Conneau et al. (2020); RetroMAE+Fine-tuning refers to the fine-tuning on the pre-trained model from RetroMAE Xiao et al. (2022). Meanwhile, RetroMAE+Unsup+Fine-tuning involves fine-tuning on a model that is trained with RetroMAE and then pre-trained on unsupervised data. The results are presented in Table 6. We can observe that RetroMAE can significantly improve the retrieval performance, and pre-training on unsupervised data can further enhance the retrieval quality of the embedding model. (More detailed results are available in Appendix C.1.)
@@ 5 Conclusion
In this paper, we introduce M3-Embedding, which substantially advances the versatility of text embeddings in terms of supporting multi-lingual retrieval, handling input of diverse granularities, and unifying different retrieval functionalities. M3-Embedding presents three technical contributions: self-knowledge distillation, efficient batching, and high-quality curation of data. The effectiveness of M3-Embedding is empirically verified, where it leads to superior performances on multi-lingual retrieval, cross-lingual retrieval, and multi-lingual long-document retrieval tasks.
Limitations
First of all, while our proposed M3-Embedding model achieves state-of-the-art performance on popular multi-lingual and cross-lingual benchmarks such as MIRACL and MKQA, it is important to acknowledge that the generalizability of our approach to diverse datasets and real-world scenarios needs to be further investigated. Different datasets may have varying characteristics and challenges that could affect the performance of our model. Secondly, while M3-Embedding is designed to process inputs of different granularities, including long documents of up to 8192 tokens, we acknowledge that processing extremely long documents could pose challenges in terms of computational resources and model efficiency. The performance of our model on very long documents or documents exceeding the specified token limit needs to be further investigated. Furthermore, we claim support for more than 100 working languages in M3-Embedding. However, the potential variations in performance across different languages are not thoroughly discussed. Further analysis and evaluation on a broader range of languages are necessary to understand the robustness and effectiveness of our model across different language families and linguistic characteristics.
Ethics Consideration
Our work proposes a new embedding model called M3-Embedding, which is distingulished for its versality in multi-linguality, multi-functionality and multi-granularity. Because our model will be publicly avaliable, it is influenced by the inherent impacts of open-source model. Moreover, we use the multilingual data including all kinds of languages in the training of M3-Embedding. However, due to the uneven distribution of training data for different languages, the model’s performance may vary across languages, which could potentially be seen as discriminatory or unfair. We ensure that our work is conformant to the ACL Ethics Policy777https://www.aclweb.org/portal/content/acl-code-ethics.
Acknowledgements
We would like to thank anonymous reviewers for their helpful feedback, and ACL 2024 and ACL Rolling Review organizers for their efforts. This research is supported by National Science and Technology Major Project (2023ZD0121504).
@@ A.1 Collected Data
| Data Source | Language | Size |
| Unsupervised Data | ||
| MTP | EN, ZH | 291.1M |
| S2ORC, Wikipeida | EN | 48.3M |
| xP3, mC4, CC-News | Multi-Lingual | 488.4M |
| NLLB, CCMatrix | Cross-Lingual | 391.3M |
| CodeSearchNet | Text-Code | 344.1K |
| Total | – | 1.2B |
| Fine-tuning Data | ||
| MS MARCO, HotpotQA, NQ, NLI, etc. | EN | 1.1M |
| DuReader, T2-Ranking, NLI-zh, etc. | ZH | 386.6K |
| MIRACL, Mr.TyDi | Multi-Lingual | 88.9K |
| MultiLongDoc | Multi-Lingual | 41.4K |
@@ B.1 Experimental Hyperparameters
We adopt a further pre-trained XLM-RoBERTa888https://huggingface.co/FacebookAI/xlm-roberta-large as the foundational model. We extend the max position to 8192 and update the model via the RetroMAE Xiao et al. (2022) method. The data comprises Pile Gao et al. (2020), Wudao Yuan et al. (2021), and mC4 Raffel et al. (2020) datasets. We sampled a total of 184 million text samples from these sources, covering 105 languages. The maximum sequence length is 8192 and the learning rate is . The batch size is set to 32 and we accumulate the gradient over 16 steps. Pre-training is conducted on 32 A100(40GB) GPUs for 20,000 steps.
@@ C.1 Additional Resutls
In this section, we present additional evaluation results on the MIRACL and MKQA benchmarks. As shown in Table 12 and 13, M3-Embedding outperforms all baselines on average.
The detailed results of ablation studies of self-knowledge distillation and multi-stage training on the MIRACL dev set are shown in Table 14 and Table 15.
@@ C.2 Different Tokenizer for BM25
Using the Analyzer from Lucene999https://github.com/apache/lucene/tree/main/lucene/analysis/common/src/java/org/apache/lucene/analysis can significantly enhance the effectiveness of BM25. Lucene analyzer includes multiple steps typically including tokenization, stemming, stopword removal, etc, achieving better results than directly using the tokenzier of XLM-RoBERTa. Additionally, it’s worth noting that the vocabulary size of the tokenizer from XLM-RoBERTa is limited, resulting in fewer unique tokens after encoding documents (for example, on the MLDR dataset, the tokenizer of XLM-RoBERTa produces 1056 unique terms per article, while Lucene’s analyzer generates 1451 unique terms, which is over 37% more and will increase retrieval latency).
@@
| Dense | 95.5 | 97.6 | 98.7 | 90.7 | 91.1 | 94.0 | 97.9 | 93.8 | 94.4 | 90.5 | 97.5 | 95.5 | 95.9 | 97.2 | 99.4 | 99.1 | 96.9 | 90.9 | 98.7 |
| Sparse | 85.6 | 92.0 | 96.7 | 81.5 | 72.1 | 87.0 | 91.5 | 73.3 | 87.1 | 84.8 | 92.4 | 91.7 | 76.9 | 85.1 | 98.1 | 95.2 | 72.9 | 69.1 | 92.9 |
| Multi-vec | 96.3 | 97.8 | 98.9 | 91.7 | 92.4 | 94.9 | 98.2 | 96.1 | 95.1 | 92.5 | 98.0 | 95.9 | 96.6 | 97.3 | 99.4 | 99.2 | 97.3 | 92.4 | 99.2 |
| Dense+Sparse | 96.2 | 98.0 | 98.9 | 92.4 | 92.5 | 95.6 | 98.3 | 94.6 | 95.6 | 92.6 | 97.5 | 95.6 | 96.6 | 97.4 | 99.1 | 99.0 | 96.8 | 91.0 | 100.0 |
| All | 96.4 | 98.0 | 98.9 | 92.1 | 92.9 | 95.6 | 98.4 | 95.6 | 95.2 | 92.5 | 98.0 | 96.0 | 96.7 | 97.2 | 99.4 | 99.2 | 97.6 | 92.3 | 99.2 |
@@
| M3-w.skd | |||||||||||||||||||
| Dense | 69.2 | 78.4 | 80.0 | 56.9 | 56.1 | 60.9 | 78.6 | 58.3 | 59.5 | 56.1 | 72.8 | 69.9 | 70.1 | 78.7 | 86.2 | 82.6 | 62.7 | 56.7 | 81.8 |
| Sparse | 53.9 | 67.1 | 68.9 | 43.8 | 38.6 | 45.1 | 65.4 | 35.3 | 48.2 | 48.9 | 56.1 | 61.5 | 44.5 | 57.9 | 79.1 | 70.9 | 36.1 | 32.5 | 70.0 |
| Multi-vec | 70.5 | 79.6 | 81.0 | 59.3 | 57.8 | 62.0 | 80.1 | 59.4 | 61.5 | 58.3 | 74.5 | 71.2 | 71.2 | 79.1 | 87.9 | 83.0 | 63.7 | 58.0 | 82.4 |
| M3-w.o.skd | |||||||||||||||||||
| Dense | 68.7 | 78.0 | 79.1 | 56.4 | 55.4 | 60.3 | 78.3 | 58.2 | 59.0 | 55.1 | 72.4 | 68.8 | 69.5 | 77.8 | 85.8 | 82.5 | 63.0 | 56.0 | 80.6 |
| Sparse | 36.7 | 48.2 | 51.9 | 24.3 | 20.3 | 26.0 | 48.6 | 16.8 | 30.1 | 32.0 | 33.0 | 43.1 | 27.2 | 45.2 | 63.6 | 52.2 | 22.6 | 16.5 | 59.2 |
| Multi-vec | 69.3 | 78.7 | 80.2 | 57.6 | 56.7 | 60.5 | 79.0 | 58.4 | 59.3 | 57.5 | 74.0 | 70.3 | 70.2 | 78.6 | 86.9 | 82.1 | 61.9 | 56.7 | 78.2 |
| Fine-tune | |||||||||||||||||||
| Dense | 60.5 | 71.0 | 72.5 | 47.6 | 46.7 | 51.8 | 72.3 | 50.9 | 48.9 | 48.9 | 65.7 | 60.5 | 60.9 | 71.9 | 81.3 | 74.7 | 54.4 | 48.7 | 60.6 |
| RetroMAE + Fine-tune | |||||||||||||||||||
| Dense | 66.1 | 75.9 | 77.9 | 54.5 | 54.0 | 58.3 | 76.6 | 55.1 | 57.0 | 53.9 | 70.1 | 66.9 | 66.9 | 74.8 | 86.1 | 79.5 | 61.9 | 52.7 | 67.5 |
| RetroMAE + Unsup + Fine-tune | |||||||||||||||||||
| Dense | 69.2 | 78.4 | 80.0 | 56.9 | 56.1 | 60.9 | 78.6 | 58.3 | 59.5 | 56.1 | 72.8 | 69.9 | 70.1 | 78.7 | 86.2 | 82.6 | 62.7 | 56.7 | 81.8 |